OpenStudy (j.maule):

A pot of stew is placed on a stove to heat. The temperature of the liquid reaches 170°F, and then the pot is taken off the burner and placed on a kitchen counter. The temperature of the air in the kitchen is 76°F. If k = 0.34, the temperature of the liquid after 7 hours will be (blank) ºF. (Hint: Use Newton’s cooling model.)

OpenStudy (j.maule):

@misty1212

OpenStudy (anonymous):

dT/dt = -k ( T - Ta )

OpenStudy (j.maule):

@jayzdd do I put that in the blank?

OpenStudy (anonymous):

thats an equation you need

OpenStudy (j.maule):

@jayzdd I don't know that equation

OpenStudy (misty1212):

HII!

OpenStudy (anonymous):

solving for T T(t) = Ta + (To - Ta) * e ^(-kt)

OpenStudy (misty1212):

\[170-76=94\] compute \[94\times e^{-.34\times 7}\] then add 76 to the result